Suppose that the production function is given as follows:

TPL = 10L + 5L2 + L3

Find the total product, Marginal product and average product when L = 5. 


1
Expert's answer
2021-04-28T20:15:04-0400

When L = 5, then:

Total product is"TP = 10\u00d75 + 5\u00d75^2 + 5^3 = 300," Marginal product is "MP = TP'(L) = 10 + 10\u00d75 + 3\u00d75^2 = 135,"

and average product is "AP = TP\/L = 10 + 5\u00d75 + 5^2 = 60."